Determine the net heat transfer through the black body disc with radius1cm and mass 6g, at 00C when transferred to a surrounding temperature of 100C.
1
Expert's answer
2020-12-17T07:25:09-0500

"r = 1 \;cm = 0.01 \;m \\

T_{env}=10 \;C = 283 \;K \\

T = 0 C = 273 \;K \\

Q = σεA(T_{env}^4 – T^4)"

"σ = 5.6704 \times 10^{-8} \; \frac{W}{m^2K^4}" Stefan Boltzmann constant

ε = 1 (for black body)

"A = πr^2 \\

Q = 5.6704 \times 10^{-8} \times 1 \times 3.14 \times (0.01)^2(283^4 – 273^4) \\

Q = 0.015 W"